A 100 N box sits on a 30 degree incline. If the static coefficient of friction is 0.1, what is the magnitude of the static frictional force acting on the box?

Select one:

a. 5.0 N

b. 8.7 N

c. 10N
Answer #1

From FBD, Using Force balance along the incline:

Ff = us*N

Ff = us*W*cos A

Using given values:

Ff = 0.1*100**cos 30 deg

Ff = 8.66 N = 8.7 N

Correct option is B.

Let me know if you've any query.